Given the Production function:
Q=100+P-0.01P2+2N-0.03N2
Determine the marginal rate of technical substitution.
1
Expert's answer
2012-07-19T07:30:34-0400
Marginal Rate of Technical Substitution is the amount by which the quantity of
one input (n) has to be reduced when one extra unit of another input (n2) is
used, so that the total output (q) remains constant:

Dn/DN2 =
MP(N2)/MP(N) = -0.03/2= -0.015 – MRTS for the input N for N2